OpenStudy (anonymous):

can anyone help? [7x/(x+3)]+[6x/(x-7)]

OpenStudy (anonymous):

\[\frac{ 7x }{ x+3 }+\frac{ 6x }{ x-7 }=\frac{ 7x(x-7)+6x(x+3) }{ (x+3)(x-7) }\]

OpenStudy (anonymous):

Expand numerator and factorise.

OpenStudy (anonymous):

take the LCD is (x+3)(x-7) and then multiply both side (x+3)(x-7)[7x/x+3 ] + (x+3)(x-7)[6x/x-7] then cross out the like tearms (x-7)7x+ (x+3)6x then multiply it

OpenStudy (anonymous):

so...its 13x^2-68x ?

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

Wait let me check.

OpenStudy (anonymous):

Incorrect @bfinn4